OpenStudy (anonymous):

Simplify m^5 N^4 --------- (M^5)^2 n (hope everyone can understand that...)

OpenStudy (anonymous):

\[\frac{m^5n^4}{(m^5)^2}\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

\[\frac{m^5n^4}{(m^5)^2n}\]

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

The power rule states that when you raise an exponent to an exponent, you multiply the exponents

OpenStudy (anonymous):

\[\frac{m^5n^4}{m^{10}n}\]

OpenStudy (anonymous):

To divide exponents, you subtract.

OpenStudy (anonymous):

\[m^{-5}n^3\]

OpenStudy (anonymous):

\[\frac{n^3}{m^5}\] \[x^{-y}=\frac{1}{x^y}\]

OpenStudy (anonymous):

how did you get ^3 from the ^4?

OpenStudy (anonymous):

\[\frac{n^4}{n^1}=n^{4-1}\]

OpenStudy (anonymous):

Subtract exponents. 4-1=3

OpenStudy (anonymous):

Okay Thank You so much ^-^